# Guiding LLMs to Generate High-Fidelity and High-Quality Counterfactual Explanations for Text Classification

### v1

@@ 5.3 More Examples in Few-shot Prompting Do Not Necessarily Improve Performance
Table 3 shows that increasing the number of shots improves the flip rate on SNLI (except for GPT-4o-mini), as well as distance and perplexity on IMDB, but has minimal effect on flip rate for IMDB. The impact of the number of shots varies by dataset and LLM type; for example, Llama-2 shows a strong decrease in distance when the number of shots increases on SNLI. On IMDB and across all other models on both datasets, differences are minimal or nonexistent. Additionally, increasing the number of shots can degrade LLM performance, especially for models with smaller context lengths like Llama-2. For instance, the flip rate on IMDB drops significantly with 5 shots, and Llama-2 cannot generate meaningful CFs with 10 shots on IMDB due to exceeding its 4096-token context limit. Therefore, we adopt a 1-shot setting for the representative results, even though n-shot settings may outperform it in certain cases.
Table 3: Influence of the number of shots (# shots) for few-shot prompting. Results for counterfactual metrics (FR: flip rate, Dis: token-level Levensthein Distance, PP: perplexity) on IMDB and SNLI across different LLMs. Bold denotes the best value across all models, and the second-best value is underlined. 10-shot prompting is not applicable (n.a.) for Llama-2 since the prompt is larger than the size of the model’s context.
@@
FR $\uparrow$ Dis $\downarrow$ PP $\downarrow$ FR $\uparrow$ Dis $\downarrow$ PP $\downarrow$
@@ 5.4 LLM-Generated CFs Are Helpful for Data Augmentation
We evaluated whether the generated CFs improve model robustness by following the experimental setup from [14]. Specifically, we augmented the training data from [14] with our generated CFs and trained a BERT classifier on the augmented data. We then evaluated the classifier on three test sets: the original test set, a CF test set created through crowdsourcing [14], and a CF test set created by the dataset authors [10]. The latter two are considered out-of-distribution data [10], where models typically struggle. As Table 4 shows, the CFs generated by our methods improve classifier accuracy on all datasets, except for GPT-4o-mini on the IMDB dataset. On the original test set, our CF-augmented data outperformed human-augmented data. However, on the more challenging human-generated test sets, LLM-generated CFs perform worse than human-augmented data but still outperform models trained without data augmentation and also outperform other counterfactual generation methods such as FLARE and MICE. These results demonstrate that our methods not only enhance CF generation but also improve downstream task performance.
@@ 5.5 LLMs Are Not Faithful to the Classifier, but Rely on Parametric Knowledge
When a classifier performs well, its decisions align with real-world logic, making it challenging to determine whether LLMs explain the classifier’s reasoning or rely on real-world knowledge that coincidentally aligns with the classifier’s decisions. Conversely, poorly performing classifiers often deviate from real-world logic. If LLMs can generate effective CFs for such classifiers, it indicates faithfulness to the classifier’s reasoning rather than external knowledge. To evaluate this faithfulness, we tested LLMs using a low-accuracy classifier. Specifically, we trained a logistic regression classifier on the IMDB dataset with reversed labels (i.e., positive swapped with negative) while keeping the test set unchanged, resulting in 17% accuracy. This setup forces LLMs to explain a classifier that operates contrary to their parametric knowledge [33]. If LLMs could successfully explain this classifier, it would demonstrate faithfulness to any classifier, regardless of performance.
Table 5: Low-accuracy classifier. Results for counterfactual metrics (FR: flip rate, Dis: token-level Levenshtein Distance, PP: perplexity) on IMDB. $\Delta$ indicates the difference compared to the same setting with the high-accuracy classifier.
FR $\uparrow$($\Delta$) Dis $\downarrow$($\Delta$) PP $\downarrow$($\Delta$) MR
@@
Table 5 shows that LLMs struggle to explain the low-accuracy classifier, with a significantly lower FR compared to the high-accuracy classifier. This suggests that LLMs heavily rely on their parametric knowledge, despite being prompted to adapt to reversed labels, aligning with prior findings in other tasks [5, 17]. GPT-4o-mini performs best, reflecting better instruction understanding, though it still leans on parametric knowledge. CGV is the most robust method, maintaining moderate flip rates on Llama-3.1 and GPT-4o-mini, while Llama-2 and CGG show sharp declines, indicating an over-reliance on parametric knowledge. Distance and perplexity also worsen, further illustrating the difficulty of explaining low-accuracy classifiers. These results emphasize the importance of evaluating LLM-based counterfactual generation on both high- and low-accuracy classifiers. While low-accuracy classifiers are unlikely to be deployed in practice, they provide a proxy for wrong classifier decisions and more importantly, allow us to evaluate the faithfulness of counterfactual explanations, i.e., whether counterfactuals align with the classifier’s true reasoning.
@@ 6 Conclusion
Our analysis uncovers a critical issue in counterfactual explanations by LLMs: LLMs are not fully faithful to the classifier, often relying more on parametric knowledge than the classifier’s actual reasoning. Developers of future LLM-based counterfactual generation methods should pay attention to this issue. We strongly recommend evaluating counterfactuals with both high- and low-accuracy classifiers to ensure faithfulness of counterfactual explanations.
@@ Appendix 0.A Example Prompts
Fig. 4 and Fig. 5 present example prompts for sentiment analysis and NLI tasks used in the CGG approach.
@@ Appendix 0.B SHAP as Feature Importance Method
The results for CGG and CGGV using SHAP [18] in Table 6 indicate that using SHAP as the XAI method to identify important words performs similarly as saliency maps, reinforcing a limitation of XAI methods: words important for the current prediction are not necessarily important for counterfactual generation.
Table 6: SHAP as feature extractor. Results for counterfactual metrics (FR: flip rate, Dis: token-level Levensthein Distance, PP: perplexity) on IMDB and SNLI across different LLMs. Bold denotes the best value across all settings.
@@
FR $\uparrow$ Dis $\downarrow$ PP $\downarrow$ FR $\uparrow$ Dis $\downarrow$ PP $\downarrow$

### v2

@@ 5.3 More Examples in Few-shot Prompting Do Not Necessarily Improve Performance
Table˜3 shows that increasing the number of shots improves the flip rate on SNLI (except for GPT-4o-mini), as well as distance and perplexity on IMDB, but has a minimal effect on the flip rate for IMDB. The impact of the number of shots varies by dataset and LLM type; for example, Llama-2 shows a strong decrease in distance when the number of shots increases on SNLI. On IMDB and across all other models on both datasets, differences are minimal or non-existent. Additionally, increasing the number of shots can degrade LLM performance, especially for models with smaller context lengths like Llama-2. For instance, the flip rate on IMDB drops significantly with 5 shots, and Llama-2 cannot generate meaningful CFs with 10 shots on IMDB due to exceeding its 4096-token context limit. Therefore, we adopt a 1-shot setting for the representative results, even though n-shot settings may outperform it in certain cases.
Table 3: Influence of the number of shots (# shots) for few-shot prompting. Results for counterfactual metrics (FR: flip rate, Dis: token-level Levensthein Distance, $P$: perplexity) on IMDB and SNLI across different LLMs. Bold denotes the best value across all models, and the second-best value is underlined. 10-shot prompting is not applicable (n.a.) for Llama-2 since the prompt is larger than the size of the model’s context.
@@
FR $\uparrow$ Dis $\downarrow$ $P$ $\downarrow$ FR $\uparrow$ Dis $\downarrow$ $P$ $\downarrow$
@@ 5.4 LLM-Generated CFs Are Helpful for Data Augmentation
We evaluated whether the generated CFs improve model robustness by following the experimental setup from [Kaushik-etal-2020-learning]. Specifically, we augmented the training data from [Kaushik-etal-2020-learning] with our generated CFs and trained a BERT classifier on the augmented data. We then evaluated the classifier on three test sets: the original test set, a CF test set created through crowdsourcing [Kaushik-etal-2020-learning], and a CF test set created by the dataset authors [gardner-etal-2020-evaluating]. The latter two are considered out-of-distribution data [gardner-etal-2020-evaluating], where models typically struggle. As Table˜4 shows, the CFs generated by our methods improve classifier accuracy on all datasets, except for GPT-4o-mini on the IMDB dataset. On the original test set, our CF-augmented data outperformed human-augmented data. However, on the more challenging human-generated test sets, LLM-generated CFs perform worse than human-augmented data but still outperform models trained without data augmentation and also outperform other counterfactual generation methods such as FLARE and MICE. These results demonstrate that our methods not only enhance CF generation but also improve downstream task performance.
@@ 5.5 LLMs Are Not Faithful to the Classifier, but Rely on Parametric Knowledge
When a classifier performs well, its decisions align with real-world logic, making it challenging to determine whether LLMs explain the classifier’s reasoning or rely on real-world knowledge that coincidentally aligns with the classifier’s decisions. Conversely, poorly performing classifiers often deviate from real-world logic. If LLMs can generate effective CFs for such classifiers, it indicates faithfulness to the classifier’s reasoning rather than external knowledge. To evaluate this faithfulness, we tested LLMs using a low-accuracy classifier. Specifically, we trained a logistic regression classifier on the IMDB dataset with reversed labels (i.e., positive swapped with negative) while keeping the test set unchanged, resulting in 17% accuracy. This setup forces LLMs to explain a classifier that operates contrary to their parametric knowledge [yu-etal-2024-revealing]. If LLMs could successfully explain this classifier, it would demonstrate faithfulness to any classifier, regardless of performance.
Table 5: Low-accuracy classifier. Results for counterfactual metrics (FR: flip rate, Dis: token-level Levenshtein Distance, $P$: perplexity) on IMDB. $\Delta$ indicates the difference compared to the same setting with the high-accuracy classifier.
FR $\uparrow$($\Delta$) Dis $\downarrow$($\Delta$) $P$ $\downarrow$($\Delta$) MR
@@
Table˜5 shows that LLMs struggle to explain the low-accuracy classifier, with a significantly lower FR compared to the high-accuracy classifier. This suggests that LLMs heavily rely on their parametric knowledge, despite being prompted to adapt to reversed labels, aligning with prior findings in other tasks [chen-etal-2022-rich, longpre-etal-2021-entity]. GPT-4o-mini performs best, reflecting better instruction understanding, though it still leans on parametric knowledge. CGV is the most robust method, maintaining moderate flip rates on Llama-3.1 and GPT-4o-mini, while Llama-2 and CGG show sharp declines, indicating an over-reliance on parametric knowledge. Distance and perplexity also worsen, further illustrating the difficulty of explaining low-accuracy classifiers. These results emphasize the importance of evaluating LLM-based counterfactual generation on both high- and low-accuracy classifiers. While low-accuracy classifiers are unlikely to be deployed in practice, they provide a proxy for wrong classifier decisions and more importantly, allow us to evaluate the faithfulness of counterfactual explanations, i.e., whether counterfactuals align with the classifier’s true reasoning.
@@ 6 Conclusion
Our analysis uncovers a critical issue in counterfactual explanations by LLMs: LLMs are not fully faithful to the classifier, often relying more on parametric knowledge than the classifier’s actual reasoning. Developers of future LLM-based counterfactual generation methods should pay attention to this issue. We strongly recommend evaluating counterfactuals with both high- and low-accuracy classifiers to ensure the faithfulness of counterfactual explanations.
@@ Appendix 0.A Example Prompts
Fig.˜4 and Fig.˜5 present example prompts for sentiment analysis and NLI tasks used in the CGG approach.
@@ Appendix 0.B SHAP as Feature Importance Method
The results for CGG and CGGV using SHAP [lundberg2017shap] in Table˜6 indicate that using SHAP as the XAI method to identify important words performs similarly as saliency maps, reinforcing a limitation of XAI methods: words important for the current prediction are not necessarily important for counterfactual generation.
Table 6: SHAP as feature extractor. Results for counterfactual metrics (FR: flip rate, Dis: token-level Levensthein Distance, $P$: perplexity) on IMDB and SNLI across different LLMs. Bold denotes the best value across all settings.
@@
FR $\uparrow$ Dis $\downarrow$ $P$ $\downarrow$ FR $\uparrow$ Dis $\downarrow$ $P$ $\downarrow$